# Efficient classical algorithms for linear optical circuits

### v1

@@ Abstract
We present efficient classical algorithms to approximate expectation values and probability amplitudes in linear optical circuits. Specifically, our classical algorithm efficiently approximates the expectation values of observables in linear optical circuits for arbitrary product input states within an additive error under a mild condition. This result suggests that certain applications of linear optical circuits relying on expectation value estimation, such as photonic variational algorithms, may face challenges in achieving quantum advantage. In addition, the (marginal) output probabilities of boson sampling with arbitrary product input states can be efficiently approximated using our algorithm, implying that boson sampling can be efficiently simulated if its output probability distribution is polynomially sparse. Moreover, our method generalizes Gurvits’s algorithm, originally designed to approximate the permanent, to also approximate the hafnian of complex symmetric matrices with an additive error. The algorithm also solves a molecular vibronic spectra problem for arbitrary product input states as precisely as boson samplers. Finally, our method extends to near-Clifford circuits, enabling the classical approximation of their expectation values of any observables and (marginal) output probabilities.
@@
Introduction.— Quantum optical systems are promising platforms to realize quantum computational advantages. In particular, linear optical circuits are experimentally feasible and can be used to build universal quantum computers using the Knill-Laflamme-Milburn scheme [1]. Also, they are considered a promising platform to demonstrate the computational advantage in the near future since the seminal work on boson sampling [2]. The boson sampling proposal [2, 3, 4] motivated various experiments to demonstrate quantum advantages [5, 6, 7, 8, 9] while numerous classical algorithms have also been developed to challenge them, leading to significant advancements in both theoretical and experimental aspects of linear optics [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
With the rapid progress in boson sampling experiments, various proposals have emerged to leverage its advantage for solving practical problems [22, 23, 24, 25, 26, 27], though their quantum advantage remains unproven. A representative example is the molecular vibronic spectra problem [22], a chemistry problem solvable with Gaussian boson sampling [3, 4], which corresponds to computing grouped output probabilities of boson sampling. While many experiments have been conducted to demonstrate this proposal [28, 29, 30], Ref. [31] has recently introduced a classical algorithm that achieves comparable performance. Its key idea is that since Gaussian boson sampling computes quantities through sampling, the accuracy is limited by the finite sample size, which leaves room for a classical algorithm to achieve the same performance. This example illustrates that errors from finite sampling must be properly accounted for in order to rigorously evaluate any potential quantum advantage in computing physical quantities via sampling.
Another notable proposal for applying boson sampling to practical problems is variational quantum computing [32, 33, 34, 35, 36]. Similarly, boson samplers play a key role in these problems by computing physical quantities through sampling, meaning that the results are again inherently approximations with additive errors due to finite sampling. This potentially allows classical algorithms to match their performance. Therefore, understanding the complexity of approximating expectation values of physical quantities in linear optical circuits is crucial for assessing the potential quantum advantage of boson sampling in practical applications.
@@
Figure 1: Summary of various classical algorithms introduced in the main text. (a) A schematic of the general expectation value of an $M$-mode linear optical circuit (Theorem 1) or an $M$-qubit near-Clifford circuit (Theorem 3). If $\hat{O}_{i}$ are photon-number projectors $|m_{i}\rangle\langle m_{i}|_{i}$, the expectation values represent marginal output probabilities of a boson sampling circuit, resulting in an efficient simulation of a sparse boson sampling (Corollary 1). (b) A probability amplitude of a linear optical circuit stands for generalized Gurvits’s algorithm (Theorem 2). When $|\phi_{i}\rangle$’s are single-photon Fock states and $|\psi_{i}\rangle$’s are squeezed vacuum states, the amplitude is proportional to the hafnian of a general complex symmetric matrix (Corollary 2). A probability amplitude is reduced to a phase shifter expectation value, Eq. (8), with $|\phi_{i}\rangle=|\psi_{i}\rangle$ and $\hat{U}\rightarrow\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ (Corollary 3).
@@
In this work, we present a classical algorithm that efficiently approximates the expectation value of operators with respect to the output states of a linear optical circuit applied to an arbitrary product input state, provided that the norm of the operator of interest is well-bounded. Our algorithm suggests that certain applications proposed to leverage boson sampling’s quantum advantage can be dequantized if the purpose of using boson sampling is to approximate physical quantities, such as cost functions in variational quantum algorithms [34, 35]. In addition, our classical algorithm enables an efficient approximation of the (marginal) output probabilities of boson sampling circuits for arbitrary input states. Thus, it proves that any boson sampling can be classically simulated if the output probability distribution is polynomially sparse [37], extending previous results that applied only to restricted cases [38, 39, 40].
@@
Furthermore, we present another classical algorithm that fully generalizes Gurvits’s algorithm [41, 2], which efficiently approximates probability amplitudes of linear optical circuits. It provides a quantum-inspired classical algorithm to approximate matrix functions, such as hafnian [42], by mapping the matrix into a Gaussian boson sampling’s output amplitude [3, 43]. Our algorithm can also efficiently approximate the expectation value of a phase shifter. The latter enables us to solve a generalized molecular vibronic spectra problem with arbitrary input states, a candidate for achieving a practical quantum advantage of boson sampling [31]. Additionally, it can be used for approximating binned probabilities of linear optical circuits with arbitrary input states, which have been proposed for verifying boson sampling experiments [44, 45, 46, 47].
@@
Finally, using a similar procedure, we extend our algorithm to near-Clifford circuits [48, 49] and show that we can efficiently estimate the expectation values of observables and any (marginal) output probabilities for these circuits within an additive error.
@@
Consider an $M$-mode linear optical circuit $\hat{U}$ and $K$ operators ${\{\hat{O}^{(k)}\}_{k=1}^{K}}$. The expectation values $\langle\psi|\hat{U}^{\dagger}\hat{O}^{(k)}\hat{U}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$.
@@
where $q(\bm{\alpha})\equiv|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}/(\pi^{l}\|\hat{O}_{A}\|_{2}^{2})$ is normalized to 1 and thus a proper probability distribution of $\bm{\alpha}$. Thus, the expression implies that by sampling $\bm{\alpha}$ from $q(\bm{\alpha})$ and averaging the random variable $X(\bm{\alpha})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})/\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})$ over the samples, we obtain an estimate of the expectation value.
We then show that the procedure can be efficiently performed and that the median-of-means estimator [54] yields an estimate with additive error $\epsilon$ with probability $1-\delta$ when $O(\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}\log\delta^{-1}/\epsilon^{2})$ number of samples is given, by showing that the variance of the random variable $X(\bm{\alpha})$ is upper-bounded by $\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}$. Finally, due to the union bound, we can approximate $K$ observables using $O(\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$ samples. Since it takes $O(M^{2})$ time for computing random variable $X(\bm{\alpha})$, the total running time is $O(M^{2}\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$.
@@
Theorem 1 has various intriguing implications. First, any applications of linear optical circuits, such as boson sampling, that rely on estimating a physical observable $\hat{O}$ can be dequantized if the target observable’s norm is properly bounded; thus, this establishes a boundary for a classically simulable regime when it comes to approximating expectation values. Furthermore, when $\hat{O}$ is an orthogonal projector, its expectation value directly corresponds to the associated output probability. Therefore, if the rank of the orthogonal projector is well-bounded, the output probability of a linear optical circuit can be efficiently approximated. A notable example is a photon-number projector in boson sampling [2]. In this case, Theorem 1 implies that the (marginal) output probabilities of boson sampling with arbitrary input states can be efficiently approximated. To the best of our knowledge, such a result has only been known for the Fock state input [2, 45, 31]. Remarkably, this finding also enables an efficient simulation of boson sampling with arbitrary input states if polynomially many outcomes dominate the output probabilities [37, 55].
@@
We emphasize that Corollary 1 is valid for a general case regardless of phase space negativity [56, 40] or stellar rank [57] of the input state, resources for quantum advantages.
@@
The probability amplitude $\langle\phi|\hat{U}|\psi\rangle$ of a linear optical circuit $\hat{U}$ with arbitrary states $|\psi\rangle$ and $|\phi\rangle$ can be efficiently approximated.
@@
While Ref. [31] proves that this is the case when $|\psi\rangle$ is a product Gaussian state or a Fock state, leading to dequantization, for more general input states, like a product squeezed Fock state, a quantum advantage may still exist since no classical algorithm is known yet. Notably, by setting $|\phi\rangle=|\psi\rangle$ and $\hat{U}\to\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ in Theorem 2, our classical algorithm can approximate Eq. (8) for any product input state:
@@
Extending to qubit circuits.— Finally, we show that our classical algorithm in Theorem 1 can be extended to near-Clifford circuits $\hat{U}_{\text{NC}}$ [48, 49], i.e., Clifford circuit $\hat{U}_{\text{C}}$+logarithmic number of $T$ gates, with arbitrary input states. Crucial properties used in Theorem 1 are that a displacement operator is transformed into another displacement operator through linear optical circuits and that any operator can be expanded in terms of displacement operators. Similar properties hold for Clifford circuits where any operators can be decomposed using Pauli operators and a Pauli operator transforms another Pauli operator under Clifford circuits. Thus, we obtain a similar result:
@@
Consider an $M$-qubit near-Clifford circuit $\hat{U}_{\text{NC}}$ with depth $O(\text{poly}(M))$ and $K$ operators ${\{\hat{O}^{(k)}\}_{k=1}^{K}}$. The expectation values $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}^{(k)}\hat{U}_{\text{NC}}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(\text{poly}(M)\max_{k}\lVert\hat{O}_{A_{k}}^{(k)}\rVert^{2}_{2}\|\hat{\rho}_{A_{k}}\|_{2}^{2}\log(K/\delta)/\epsilon^{2})$.
@@
Thus, by sampling $\bm{a}$ from $q(\bm{a})\equiv|\chi_{\hat{O}_{A}}(\bm{a})|^{2}/(2^{l}\|\hat{O}_{A}\|_{2}^{2})$ and averaging the random variable $X(\bm{a})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})/\chi^{*}_{\hat{O}_{A}}(\bm{a})$ over the samples, we obtain an estimate of the expectation value. By showing that the variance of the random variable $X(\bm{a})$ is bounded by $\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}$, the median-of-means estimator together with the union bound gives the desired error presented in the theorem.
@@
Theorem 3 implies that the expectation value of observables $\hat{O}$ with $\lVert\hat{O}\rVert_{2}^{2}=O(\text{poly}(M))$ and marginal probabilities of a near-Clifford circuit with an arbitrary input state can be efficiently approximated. Thus, for approximating the expectation values of near-Clifford circuits, the magic of input states does not increase the complexity [63]. Hence, our result is a generalization of previous results: the expectation values of a near-Clifford circuit for $|\psi\rangle=|0^{M}\rangle$ [48, 49] and (marginal) output probabilities of a Clifford circuit [64, 55].

### v2

@@ Abstract
We present efficient classical algorithms to approximate expectation values and probability amplitudes in linear optical circuits. Specifically, our algorithm efficiently approximates the expectation values of observables in linear optical circuits for arbitrary product input states within an additive error under a mild condition. This result suggests that certain applications of linear optical circuits relying on expectation value estimation, such as photonic variational algorithms, may face challenges in achieving quantum advantage. In addition, the (marginal) output probabilities of boson sampling with arbitrary product input states can be efficiently approximated using our algorithm, implying that boson sampling can be efficiently simulated if its output probability distribution is polynomially sparse. Our method generalizes Gurvits’s algorithm, originally designed to approximate the permanent, to also approximate the hafnian of complex symmetric matrices with an additive error. The algorithm also solves a molecular vibronic spectra problem for arbitrary product input states as precisely as boson samplers. Finally, our method extends to near-Clifford circuits, enabling the classical approximation of their expectation values of any observables and (marginal) output probabilities.
@@
Introduction.— Linear optical system has been considered as a promising platform for universal quantum computing [1] and also for demonstrating the quantum computational advantage in the near future since the seminal work on boson sampling [2]. The boson sampling proposal [2, 3, 4] motivated various experiments to demonstrate the quantum advantage [5, 6, 7, 8, 9] while numerous classical algorithms have also been developed to challenge them, leading to significant advancements in understanding the computational complexity of linear optical circuits [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
Lately, the main direction of this research has been shifting toward achieving practical quantum advantage: finding a useful problem that cannot be efficiently solved by classical computers but can be addressed by exploiting linear optical circuits [23, 24, 25, 26, 27, 28, 29, 30, 31]. One promising direction of exploiting sampling devices involves problems where the goal is to compute physical quantities through sampling, such as in Monte Carlo methods. Especially, many such problems can be formulated as estimating expectation values of observables, often referred to as the quantum mean value problem (e.g., Ref. [32]). Although various proposals have been suggested in this direction with boson sampling [23, 28, 29, 30, 31], no provable advantage has yet been demonstrated; thus, these approaches cannot rule out the existence of efficient classical algorithms. For example, the molecular vibronic spectra problem, which has been considered a promising potential application, can be efficiently solved using Gaussian boson sampling [23, 3, 4], and several experiments have been conducted to validate this proposal [33, 34, 35]. However, a recent result shows that a classical algorithm can achieve comparable performance [36], ruling out the practical advantage of boson sampling for this problem.
The biggest challenge in leveraging the quantum advantage of boson sampling for solving the associated quantum mean value problem is the limited accuracy due to the finite sample size, as in the above example. This limitation ultimately leaves room for a classical algorithm to achieve the same performance. Thus, identifying the classical limit on the achievable accuracy for the quantum mean value problem is crucial to rigorously establish a practical quantum advantage. Although classical algorithms for the quantum mean value problem have been developed in various contexts [32, 37, 38, 39, 40, 41], no such algorithm has yet been studied for linear optical circuits, to the best of our knowledge.
@@
Figure 1: Summary of classical algorithms introduced in the main text. (a) A schematic of the general expectation value of an $M$-mode linear optical circuit (Theorem 1) or an $M$-qubit near-Clifford circuit (Theorem 3). If $\hat{O}_{i}$ are photon-number projectors $|m_{i}\rangle\langle m_{i}|_{i}$, the expectation values represent marginal output probabilities of a boson sampling circuit, resulting in an efficient simulation of a sparse boson sampling (Corollary 1). (b) A probability amplitude of a linear optical circuit stands for generalized Gurvits’s algorithm (Theorem 2). When $|\phi_{i}\rangle$’s are single-photon Fock states and $|\psi_{i}\rangle$’s are squeezed vacuum states, the amplitude is proportional to the hafnian of a general complex symmetric matrix (Corollary 2). A probability amplitude is reduced to a phase shifter expectation value, Eq. (8), with $|\phi_{i}\rangle=|\psi_{i}\rangle$ and $\hat{U}\rightarrow\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ (Corollary 3).
@@
In this work, we present an efficient classical algorithm for approximating the expectation values in linear optical circuits. Our algorithm works for arbitrary product input states when the norm of the operator of interest is well-bounded, dequantizing a wide range of potential applications, such as computing cost functions in variational quantum algorithms [42, 43]. In addition, the algorithm enables an efficient approximation of the (marginal) output probabilities of boson sampling circuits for arbitrary input states. Thus, it proves that any boson sampling can be classically simulated if the output probability distribution is polynomially sparse [44], extending previous results that applied only to restricted cases [45, 46, 47].
@@
We further extend our result by identifying cases in which it remains valid even for operators with unbounded norm. One such case is the phase operator in the generalized molecular vibronic spectra problem, which was proposed as a candidate achieving practical quantum advantage [36]. We fully solve this problem by providing a classically efficient algorithm for arbitrary input states. Additionally, it can be used for approximating binned probabilities of linear optical circuits with arbitrary input states, which have been proposed for verifying boson sampling experiments [48, 49, 50, 51]. In the language of computational complexity, our work generalizes Gurvits’s algorithm [52, 2]. It provides a quantum-inspired classical algorithm to approximate matrix functions, such as hafnian [53], by mapping the matrix function of interest to a Gaussian boson sampling’s output amplitude [3, 54].
@@
Finally, using a similar procedure, we apply our method to near-Clifford circuits [55, 56] and show that we can efficiently estimate the expectation values of observables and any (marginal) output probabilities for these circuits within an additive error.
@@
Consider an $M$-mode linear optical circuit $\hat{U}$ and an operator $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}\hat{O}\hat{U}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}\log(1/\delta)/\epsilon^{2})$.
@@
where $q(\bm{\alpha})\equiv|\chi_{\hat{O}_{A}}(\bm{\alpha})|^{2}/(\pi^{l}\|\hat{O}_{A}\|_{2}^{2})$ is normalized to 1 and thus a proper probability distribution of $\bm{\alpha}$. Thus, the expression implies that by sampling $\bm{\alpha}$ from $q(\bm{\alpha})$ and averaging the random variable $X(\bm{\alpha})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}^{*}(\bm{\alpha})/\chi^{*}_{\hat{O}_{A}}(\bm{\alpha})$ over the samples, we obtain an estimate of the expectation value. Thanks to the tensor-product structure of the state $\hat{\rho}_{A}$ and the operator $\hat{O}_{A}$, this process can be performed efficiently [61, 62, 63].
We then show that the median-of-means estimator [64] yields an estimate with additive error $\epsilon$ with probability $1-\delta$ when $O(\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}\log\delta^{-1}/\epsilon^{2})$ number of samples is given, by showing that the variance of the random variable $X(\bm{\alpha})$ is upper-bounded by $\|\hat{O}_{A}\|_{2}^{2}\|\hat{\rho}_{A}\|_{2}^{2}$. Note that the median-of-means estimator for complex random variables can be constructed by combining the real and imaginary parts estimators [57]. Since it takes $O(M^{2})$ time for computing random variable $X(\bm{\alpha})$, the total running time is $O(M^{2}\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}\log(1/\delta)/\epsilon^{2})$. ∎
@@
Theorem 1 has various intriguing implications. First, any applications of linear optical circuits, such as boson sampling, that rely on estimating a physical observable $\hat{O}$ can be dequantized if the target observable’s norm is properly bounded; thus, this establishes a boundary for a classically simulable regime when it comes to approximating expectation values. Furthermore, when $\hat{O}$ is an orthogonal projector, its expectation value directly corresponds to the associated output probability. Therefore, if the rank of the orthogonal projector is well-bounded, the output probability of a linear optical circuit can be efficiently approximated. A notable example is a photon-number projector in boson sampling [2]. In this case, Theorem 1 implies that the (marginal) output probabilities of boson sampling with arbitrary input states can be efficiently approximated. To the best of our knowledge, such a result has only been known for the Fock state input [2, 49, 36]. Remarkably, this finding also enables an efficient simulation of boson sampling with arbitrary input states if polynomially many outcomes dominate the output probabilities [65, 44, 66].
@@
In addition, we note that quantum sampling with a peaked output probability distribution, i.e., one that has an outcome with an inverse-polynomially large probability, has drawn attention for a verifiable quantum advantage of sampling [67, 68] because the peak can serve as a verification. However, our result, together with the algorithm in Refs. [65, 44], indicates that the peak of the output probability distribution in boson sampling with any input states can be efficiently identified using classical computers. We emphasize that Corollary 1 is valid for a general case regardless of phase space negativity [69, 47] or stellar rank [70] of the input state, resources for quantum advantages.
@@
The probability amplitude $\langle\phi|\hat{U}|\psi\rangle$ of a linear optical circuit $\hat{U}$ with arbitrary states $|\psi\rangle$ and $|\phi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(M^{2}\log\delta^{-1}/\epsilon^{2})$.
@@
While Ref. [36] proves that this is the case when $|\psi\rangle$ is a product Gaussian state or a Fock state, leading to dequantization, for more general input states, like a product squeezed Fock state, a quantum advantage may still exist since no classical algorithm is known yet. This possibility is particularly important because such an extension has recently been experimentally implemented [35]. Notably, by setting $|\phi\rangle=|\psi\rangle$ and $\hat{U}\to\hat{U}^{\dagger}e^{i\hat{\bm{n}}\cdot\bm{\phi}}\hat{U}$ in Theorem 2, our classical algorithm can approximate Eq. (8) for any product input state:
@@
Extending to qubit circuits.— Finally, we show that our classical algorithm in Theorem 1 can be extended to near-Clifford circuits $\hat{U}_{\text{NC}}$ [55, 56], i.e., Clifford circuit $\hat{U}_{\text{C}}$+logarithmic number of $T$ gates, with arbitrary input states. Note that exact simulation of such a circuit is hard, similar to the case of boson sampling [75]. Crucial properties used in Theorem 1 are that a displacement operator is transformed into another displacement operator through linear optical circuits and that any operator can be expanded in terms of displacement operators. Similar properties hold for Clifford circuits where any operators can be decomposed using Pauli operators and a Pauli operator transforms another Pauli operator under Clifford circuits. Thus, we obtain a similar result:
@@
Consider an $M$-qubit near-Clifford circuit $\hat{U}_{\text{NC}}$ with depth $O(\text{poly}(M))$ and an operators $\hat{O}$. The expectation values $\langle\psi|\hat{U}^{\dagger}_{\text{NC}}\hat{O}\hat{U}_{\text{NC}}|\psi\rangle$ can be approximated within additive error $\epsilon$ with probability $1-\delta$ in running time $O(\text{poly}(M)\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}\log(1/\delta)/\epsilon^{2})$.
@@
Thus, by sampling $\bm{a}$ from $q(\bm{a})\equiv|\chi_{\hat{O}_{A}}(\bm{a})|^{2}/(2^{l}\|\hat{O}_{A}\|_{2}^{2})$ and averaging the random variable $X(\bm{a})\equiv\|\hat{O}_{A}\|_{2}^{2}\chi_{\hat{\rho}_{A}}(\bm{a})/\chi^{*}_{\hat{O}_{A}}(\bm{a})$ over the samples, we obtain an estimate of the expectation value. By showing that the variance of the random variable $X(\bm{a})$ is bounded by $\lVert\hat{O}_{A}\rVert^{2}_{2}\|\hat{\rho}_{A}\|_{2}^{2}$, the median-of-means estimator gives the desired error presented in the theorem.
@@
Theorem 3 implies that the expectation value of observables $\hat{O}$ with $\lVert\hat{O}\rVert_{2}^{2}=O(\text{poly}(M))$ and marginal probabilities of a near-Clifford circuit with an arbitrary input state can be efficiently approximated. Thus, for approximating the expectation values of near-Clifford circuits, the magic of input states does not increase the complexity [76]. Hence, our result is a generalization of previous results: the expectation values of a near-Clifford circuit for $|\psi\rangle=|0^{M}\rangle$ or for Pauli observables [55, 56] and (marginal) output probabilities of a Clifford circuit [77, 66].